OpenStudy (yacoub1993):

The French Club is sponsoring a bake sale. If their goal is to raise at least $140, how many pastries must they sell at $3.50 each in order to meet that goal? Write and solve an inequality.

OpenStudy (anonymous):

140 divided by 3.50

OpenStudy (anonymous):

140/3.50 = 40 pastried

OpenStudy (anonymous):

pastries*

OpenStudy (anonymous):

Since they want to make at least $140, that amount is the minimum and any amount greater than $140 is also an acceptable profit. So, the sign of the inequality will be a greater than or equal to sign. The inequality set-up will be: 140>==3.50x Where x represents the number of pastries that need to be sold. To solve this we simply need to divide both sides by $3.50 to get x by itself. When we do this we find that in order to make $140, the french club must sell at least 40 pastries.

OpenStudy (anonymous):

do you need only the direct answer or you need to come up with an equation?

OpenStudy (yacoub1993):

\[A. 3.50p \ge140;p \ge 136.5\] \[B. 3.50p \ge 140; p \ge 490\] \[C. 140p \ge 3.50; p \ge 40\] \[D. 3.50p \ge140; p \ge 40\]

OpenStudy (anonymous):

C

OpenStudy (yacoub1993):

thnxs